OpenStudy (tinney12lexi):

y=x^3-5x^2+9x-45 you have to find all the zero's.

OpenStudy (anonymous):

Try dividing x^3-5x^2+9x-45 by x-5

OpenStudy (tinney12lexi):

ok i will try that :)

OpenStudy (tinney12lexi):

it didnt work :(

OpenStudy (tinney12lexi):

@ missiey, @graveboat789, @Spring98

OpenStudy (tinney12lexi):

@gravyboat789

OpenStudy (spring98):

I'm sorry but I don't know this..

OpenStudy (tinney12lexi):

ok thanks :)

OpenStudy (anonymous):

\[\frac{x^3-5 x^2+9 x-45}{x-5}=x^2+9 \]

OpenStudy (tinney12lexi):

hey i foud the other zero; they are 5,-3i,3i....(:

OpenStudy (tinney12lexi):

*found